# A Survey of Route Recommendations: Methods, Applications, and Opportunities

### 01 Mar 2024 v1

@@ Abstract
Nowadays, with advanced information technologies deployed citywide, large data volumes and powerful computational resources are intelligentizing modern city development. As an important part of intelligent transportation, route recommendation and its applications are widely used, directly influencing citizens’ travel habits. Developing smart and efficient travel routes based on big data (possibly multi-modal) has become a central challenge in route recommendation research. Our survey offers a comprehensive review of route recommendation work based on urban computing. It is organized by the following three parts: 1) Methodology-wise. We categorize a large volume of traditional machine learning and modern deep learning methods. Also, we discuss their historical relations and reveal the edge-cutting progress. 2) Application-wise. We present numerous novel applications related to route commendation within urban computing scenarios. 3) We discuss current problems and challenges and envision several promising research directions. We believe that this survey can help relevant researchers quickly familiarize themselves with the current state of route recommendation research and then direct them to future research trends.
@@ A Survey And Simulation Study
Route recommendation or planning is key to many route-based applications such as navigation, deliveries, and travel planning (13). The main goal is to design routes that can be realized on a given transportation network with certain travel requirements. This has long been a very active research area. Traditional search-based algorithms (e.g. Dijkstra and tree-based searches) (14) convert a given physical space into an abstracted graph structure, where nodes represent locations and edges connections. The edges are further defined with weights that reflect certain travel costs. Then, given a set of nodes and some travel constraints, a search algorithm returns desired routes that can traverse through the queried nodes. While such exact search algorithms can provide very optimal solutions, they often do not scale up well. Therefore, heuristic search algorithms (e.g. A*(15), LPA* (16)) and machine learning methods based on approximate optimization have emerged. The main advantage is that they typically avoid processing global or exact information, and thus end up with an acceptable sub-optimal solution but with a significant improvement in search efficiency. More recently, route recommendation algorithms have experienced a revolution brought by deep learning. With the advances in data storage and computing technologies, deep learning (17) prevails and has demonstrated much stronger data representation and knowledge extraction abilities. It enables dealing with more complex route recommendation problems that are large-scale, multi-objective, and personalized-oriented with multi-source, multi-modal contextual information (18). Therefore, how the traditional machine learning approaches and the modern methodologies are inherently connected and improved, and what novel applications have been conceived need a thorough review from new perspectives.
@@
Our survey provides a comprehensive review of route recommendation methods and applications based on urban computing scenarios (19). We survey existing work from both traditional machine learning and modern deep learning perspectives. Our contributions are threefold:
@@
The structure of this survey is organized below. Section 2 gives a brief introduction of the background, problem definitions, and evaluation metrics for route recommendation. Section 3 overviews traditional machine learning route recommendation methods. Section 4 reviews and classifies modern route recommendation works based on deep learning. Following that, Section 5 demonstrates different categories of route recommendation applications. And finally, Section6 discusses current open problems and points out the future research directions.
@@ 2.2 Evaluation Metrics
3 Machine Learning Methods for Route Recommendation
We start with reviewing traditional route recommendation algorithms and broadly name them as machine learning (ML) methods, in contrast to modern deep learning methods. Typically, ML-based methods aim to find optimal routes according to a given objective (often single-objective), such as minimum distance, time, or cost. These algorithms, methods, or models can be roughly categorized into five groups: search-based, probability-based, biomimetic-based, clustering-based, and constraint-based. Many of these methods originated from the field of robotics, where automatic pathfinding tasks are crucial. Later, they were adapted and applied to various other domains such as urban road finding and navigation. In the following subsections, we will discuss the representative methods and their extensions for each group.
@@ 3.1 Search-based Methods
Search-based methods work by exploring all possible paths from one node to the other, and meanwhile, filtering out the paths that fail to satisfy the given constraints such as a maximum distance or cost. A search tree can often be built to construct the search space, and many search algorithms may apply, such as recursive search, dynamic programming, branch and bound method, etc. Search-based algorithms are usually simple and optimal but can also be inefficient. They are not suited to deal with large networks with dynamic changes.
@@
The rapid expansion of cities and urban road networks imposes an increasing demand for real-time computing. Therefore, efficient storage and computation methods for live searches are needed. Sanders and Schultes (26; 27) proposed the Highway Hierarchies (HH) algorithm optimized for practical applications. The main contribution is that HH reduces the size of network data when preprocessing directed graph structures, thus saving storage and improving computation speed. Geisberger et al. (28) introduced the Contraction Hierarchies (CH) algorithm, which searches and recommends the fastest path based on real-time travel time using heuristic functions. The system developed by Delling et al. (29) can support multiple types of optimization objectives for route recommendations in larger map structures, and their methods using overlay topology and parallel computing methods can speed up real-time searches and responses.
@@ 3.2 Probability-based Methods
Generally speaking, large graphs with complete sets of nodes and edges can still be hard to search, even enhanced with heuristics or pruning techniques. Therefore, sampling-based methods have been proposed, which essentially trade off search completeness for probabilistic completeness. In other words, graphs or routes can be sampled in part for construction. For example, Lavalle (43) developed a Probabilistic Road Map (PRM), which constructs a path network in a complex environment by random sampling (transforms the continuous space into a discrete one) and then performs path planning on the sampled network to improve search efficiency in high-dimensional spaces. Besides, the following works, falling in the RRT family, are methods that sample paths. First, Lavalle and Kuffner (44) proposed the Rapidly-exploring Random Tree (RRT) algorithm, which is a single-query algorithm that quickly finds a feasible path from the start point to the end point by random exploring. Its search process resembles a tree growing and expanding in different directions. Kuffner and LaValle (45) later introduced the RRT-Connect algorithm, which can be considered as a bidirectional expansion of two RRT trees. The two trees grow from the start/end point respectively, and they together generate a path when they connect. Karaman and Frazzoli (46) proposed RRT*, which is an asymptotically optimal algorithm. RRT* works similarly to RRT, except that RRT* has a reconnection process when selecting a parent node for a newly added node $v_{new}$. Instead of using the original parent node that generates $v_{new}$, RRT* selects a new parent $v_{pa}$ within a neighbor of $v_{new}$ that has the smallest path cost from $v_{pa}$ to $v_{new}$.
@@
3.3 Biomimetic Methods
Biological organisms have inspired many artificial intelligence applications including path planning. Among them, one of the simplest forms of life is called physarum polycephalum, a slime mold that belongs to the amoeba species in the protozoa. This single-celled organism exhibits an interesting phenomenon: on a flat culture medium, its cellular growth path can provide valuable insights for the design of planar transportation networks (116). Therefore, methods that observe, summarize, and apply the rules of biological behaviors in reality to route recommendation are collectively called biomimetic methods.
@@ 3.4 Clustering-based Methods
Clustering techniques are common useful tools for data dimension reduction and pattern recognition. In our route recommendation setting, clustering is usually used for finding similar route preferences among different users and/or for detecting similar road network structures and traffic flow patterns. By applying appropriate clustering techniques, the original problems can be solved structurally on a smaller scale, thus improving the search efficiency.
@@ 3.5 Constraint-based Methods
Constraint-based methods explicitly take into account the route requirements or preferences of users and model them as constraints in the optimization procedure.
@@ 4 Deep learning for Route Recommendation
Deep learning has become the mainstream for route recommendation methods in recent years. Though neural network models have been researched for decades, they are gradually becoming feasible for solving practical problems lately. This mainly benefits from the rapid development of hardware such as CPU, GPU, and data storage, which allow faster computation with large volumes of data. Compared to traditional machine learning methods, neural networks, especially deep models with delicate structural designs, have huge potential to discover very complex relationships among data. So in this section, we will see why and how deep learning methods prevail in route recommendation research.
@@ 4.1 Why Deep Learning for Route Recommendation?
In the following, we will present deep route recommendation works and categorize them into five types: hybrid models, sequence-based models, graph neural network models, multi-modal models, and deep reinforcement learning models. Hybrid methods usually extend traditional machine learning methods (such as A*) with neural networks; sequence-based methods mainly model the temporal relations with recurrent neural networks (RNNs); graph neural networks (GNNs) further model spatio-temporal relations; multi-modal methods can process more complex data coming from different modalities; and lastly, reinforcement learning methods explicitly focus on making decisions for route recommendations.
@@ 4.2 Hybrid Approaches
Recall that A* aims to model a cost function $F(x)=g(x)+h(x)$, where $g(x)$ is a base cost function for node $v_{x}$ and $h(x)$ is a heuristic function. The core of A* is to design a good $h(x)$ and estimate both $g(x)$ and $h(x)$ accurately. For example, an early attempt was (75; 81), which proposed using neural networks to automatically learn the cost functions of A* for personalized route recommendation tasks. Wu et al. (76) extended A* using neural networks to model complex traffic information and applied it to the recommendation of fast-traveling routes. Yang et al.(79) proposed a new method named Noah, which combines A* with a graph neural network to calculate an approximate graph editing distance more efficiently and intelligently. Agostinelli et al. (80) proposed and studied batch versions of the A* algorithm. They applied approximate admissible transformations to the heuristic functions that are parameterized by deep neural networks. They showed that these heuristic functions could find optimal or bounded suboptimal solutions. Liu et al. (73) designed a fuel-efficient route recommendation model for long-distance driving based on historical trajectories. They combined neural networks with heuristic algorithms to automatically learn the optimal solutions of A*. Lately, Wang et al. (74) proposed ASNN-FRR, which extracts features from trajectory data with urban road information using AGCRN (117) modules and estimates the values of $g(x)$ and $h(x)$ through a multi-layer perceptron (MLP) network. Besides A*, some works also adopted the classic Dijkstra algorithm in modern models. Jain et al. (78) proposed an inductive algorithm and a generative model called NEUROMLR for robust and reliable route recommendations. They convert the original computational task to the problem of finding the shortest path in a weighted graph and solve it by Dijkstra’s Algorithm, which has guaranteed optimality and accessibility. Liu et al.(13) utilized a deep inverse reinforcement learning (IRL) method to recommend delivery routes. They used the Dijkstra algorithm instead of value iteration to determine the current best strategy and calculate the gradients of the IRL model. Other work such as Huang et al. (77) proposed a multi-task route recommendation framework in contrast to single-planning methods that only work for specific tasks. They used deep learning techniques to extract path representations and employed beam search algorithms to provide recommendations for multiple related tasks.
@@ 4.3 Sequence-based Approaches
The following examples are RNN-based methods for route recommendation. An early work (85) used RNN to extract feature representations of different POI nodes and added maximum difference constraints to recommend a path with diverse scenes for users. Wang et al. (75; 81) proposed a personalized route recommendation method, which uses GRU segments to extract features from numerous user trajectories and employs the attention mechanism to increase the prediction accuracy. Similarly, Wen et al. (82) used GRUs to extract the temporal correlations in trajectory data, which helped to improve route recommendations for cargo deliveries. Huang et al. (77) proposed a multi-task prediction framework for route recommendation, which uses LSTM to fuse POI features, user features, and current path features, and applies a greedy network search algorithm to output results for multiple tasks. Hahn et al. (83) aimed to monitor collision information on roads and then recommend safer routes. They extracted features of collision accidents using LSTM and used them to predict the next step of moving positions. Liu et al. (73) combined a bidirectional GRU and attention mechanism to estimate the fuel consumption in long-distance driving routes and recommended energy-efficient routes for drivers. Fu and Lee (84) designed a progressive route recommendation method based on generative networks, which extracts knowledge from the selection of the next node using LSTM. Wang et al. (74) introduced a travel route recommendation algorithm using AGCRN modules (117), which combines an RNN and a graph convolutional neural network (GCN) to capture spatial and temporal correlations in a fine-grained dimensional space.
@@
Besides RNNs which are autoregressive models, the Transformers (121) are attention-based models that can process much longer sequences without suffering from the forgetting problem. Therefore, many studies began to explore route recommendation methods based on Transformer. For instance, Bhumika and Debasis Das (87) introduced a multi-task learning framework, which uses Transformers to extract spatial, temporal, and semantic correlations, and adopted epsilon constraint techniques to optimize routes that satisfy multiple objective functions. Wang et al. (88) studied long-distance travel trajectories, from which they extracted fuel-consumption factors using Transformer. Further combining a genetic optimization algorithm, they developed a model that could recommend more fuel-economic routes. Wang et al. (89) modeled route recommendation as a process of inferring a high-sampling rate fine trajectory from a low-sampling rate coarse trajectory composed of origin-destinations and way-points. Their proposed Seq2Seq model using multi-head self-attention mechanisms can automatically adjust the sampling weights to capture more accurate spatio-temporal correlation. Wang et al. (90) proposed a new de-biased representation learning method named TripRec to implement the sequence generation method from Query to Trip. Based on the encoded representations, they further used Transformer for decoding and generated final routes. Ngai Lam Ho and Kwan Hui Lim (91) proposed POIBERT, which uses the BERT language model to recommend personalized itineraries based on users’ preferences and their past POI selections.

### 06 Apr 2024 v2

@@ Abstract
Nowadays, with advanced information technologies deployed citywide, large data volumes and powerful computational resources are intelligentizing modern city development. As an important part of intelligent transportation, route recommendation and its applications are widely used, directly influencing citizens’ travel habits. Developing smart and efficient travel routes based on big data (possibly multi-modal) has become a central challenge in route recommendation research. Our survey offers a comprehensive review of route recommendation work based on urban computing. It is organized by the following three parts: 1) Methodology-wise. We categorize a large volume of classic methods and modern deep learning methods. Also, we discuss their historical relations and reveal the edge-cutting progress. 2) Application-wise. We present numerous novel applications related to route commendation within urban computing scenarios. 3) We discuss current problems and challenges and envision several promising research directions. We believe that this survey can help relevant researchers quickly familiarize themselves with the current state of route recommendation research and then direct them to future research trends.
@@ A Survey And Simulation Study
Route recommendation or planning is key to many route-based applications such as navigation, deliveries, and travel planning (13). The main goal is to design routes that can be realized on a given transportation network with certain travel requirements. This has long been a very active research area. Traditional search-based algorithms (e.g. Dijkstra and tree-based searches) (14) convert a given physical space into an abstracted graph structure, where nodes represent locations and edges connections. The edges are further defined with weights that reflect certain travel costs. Then, given a set of nodes and some travel constraints, a search algorithm returns desired routes that can traverse through the queried nodes. While such exact search algorithms can provide very optimal solutions, they often do not scale up well. Therefore, heuristic search algorithms (e.g. A*(15), LPA* (16)) and machine learning methods based on approximate optimization have emerged. The main advantage is that they typically avoid processing global or exact information, and thus end up with an acceptable sub-optimal solution but with a significant improvement in search efficiency. More recently, route recommendation algorithms have experienced a revolution brought by deep learning. With the advances in data storage and computing technologies, deep learning (17) prevails and has demonstrated much stronger data representation and knowledge extraction abilities. It enables dealing with more complex route recommendation problems that are large-scale, multi-objective, and personalized-oriented with multi-source, multi-modal contextual information (18). Therefore, how the classic methods and the modern methodologies are inherently connected and improved, and what novel applications have been conceived need a thorough review from new perspectives.
@@
Our survey provides a comprehensive review of route recommendation methods and applications based on urban computing scenarios (19). We survey existing work from both classic methods and modern deep learning perspectives. Our contributions are threefold:
@@
The structure of this survey is organized below. Section 2 gives a brief introduction of the background, problem definitions, evaluation metrics, and datasets for route recommendation. Section 3 overviews classic methods route recommendation methods. Section 4 reviews and classifies modern route recommendation works based on deep learning. Following that, Section 5 demonstrates different categories of route recommendation applications. And finally, Section6 discusses current open problems and points out the future research directions.
@@ 2.2 Evaluation Metrics
Apart from the above metrics, in real-world application scenarios such as taxi revenue, travel route recommendations, etc., there will be more particular ways to evaluate results, and those will be discussed inline with certain applications in Section  5.
Table 2: Route recommendation datasets.
Domain Dateset & Link About
Route Networks
City and Country Maps
Intersection IDs, road structure, POI, subway stations, etc.
Mario A. Nascimento (21)
New York Road Network (22)
Security-related
Chicago Crime Record (23)
Criminal records, including regions, types of crimes,
Philadelphia Crime Record (23)
pursuit situations, etc. Meteorological Agency data mainly
Beijing Crime Hot Spots (23)
include information such as weather, disaster types,
Japan Meteorological Agency (24)
longitude, and latitude.
Tourism-related
YFCC100M (25)
New York Road POI (22)
POI Pictures and corresponding text comments, geographical
NYC TCL record (26)
location information of scenic spots.
MATSim (27)
Indoor-related
Station plans from Swiss federal railway SBB (28)
Indoor-related datasets, including information such as store IDs, floors, and locations. They contain information such
Indoor Venue Keyword Dataset (29)
as brand names, store sizes, types, etc.
Constraint-related
New York transportation network (30)
New York City bus and subway stations, public transportation routes, and schedules.
Economics-related
Food Pick-up and Delivery Data (31)
Taxi-BJ
This mainly includes information such as taxi ID, driving
Taxi-NYC (2)
trajectory, passenger status, and speed.
DiDi
Porto Taxi (32)
2.3 Datasets
Lastly, we present common datasets relevant to route recommendation. Since route recommendation tasks mainly use trajectory data and road network information, additional auxiliary data such as passenger status and scenic spots are often included as well. As some datasets are essentially private and require access permissions, we here mainly list publicly accessible datasets, shown in  Table 2. The list has included map data, trajectory data, POI and attraction data, vehicle speed data, etc. The categories are chosen consistent to the application scenarios to be presented in Section 5.
In the list, road network data include the basic construction of urban maps and the calibration of some POIs. For example, the Mario dataset covers the intersection IDs and latitudes within a city, with the edge data representing the road structure and connectivity of the city. POI data include information such as the location of well-known buildings, bus stops, and subway stations in a city. Security-related data contain criminal records and natural disaster records. Among them, criminal records include different regions, types of crimes, pursuit situations, etc. Natural disaster data mainly include information such as weather, disaster types, longitude, and latitude. Tourism-related data mainly include pictures and corresponding text comments of scenic spots, as well as geographical location information of scenic spots. Indoor-related data cover information such as store IDs, floors, and locations. They can further contain brand names, store sizes, types, etc. Constraint-related data offer information on New York City bus and subway stations, public transportation routes, and schedules. Economics-related data mainly focus on taxis. This includes information such as taxi ID, driving trajectory, passenger status, and speed. It can reflect the income level of taxis over a while and better fit the revenue objective function to calculate route recommendation results.
@@ 2.3 Datasets
3 Classic Methods for Route Recommendation
We start with reviewing classic route recommendation algorithms. This section mainly covers non-deep-learning methods which usually have a long history, so we name them as Classic Methods (CM). Although classic, they still lay the foundation for many modern methods. Typically, CM aims to find optimal routes according to a given objective (often single-objective), such as minimum distance, time, or cost. These algorithms, methods, or models can be roughly categorized into five groups: search-based, probability-based, biomimetic-based, clustering-based, and constraint-based. Many of these methods originated from the field of robotics, where automatic pathfinding tasks are crucial. Later, they were adapted and applied to various other domains such as urban road finding and navigation. In the following subsections, we will discuss the representative methods and their extensions for each group.
@@ 3.1 Search-based Methods
Search-based methods work by exploring all possible paths from one node to the other, and meanwhile, filtering out the paths that fail to satisfy the given constraints such as a maximum distance or cost. A search tree can often be built to construct the search space, and many search algorithms may apply, such as recursive search, dynamic programming, branch and bound method, etc. Search-based algorithms are usually simple and optimal but can also be inefficient. They are not suited to deal with large networks with dynamic changes. An example process of the search-based method is shown in Figure 3.
Figure 3: An example process of search-based methods. Featured by strategically searching for available paths that connect the starting node and the end node, and finally returning the computed optimal routes.
@@
The rapid expansion of cities and urban road networks imposes an increasing demand for real-time computing. Therefore, efficient storage and computation methods for live searches are needed. Sanders and Schultes (38; 39) proposed the Highway Hierarchies (HH) algorithm optimized for practical applications. The main contribution is that HH reduces the size of network data when preprocessing directed graph structures, thus saving storage and improving computation speed. Geisberger et al. (40) introduced the Contraction Hierarchies (CH) algorithm, which searches and recommends the fastest path based on real-time travel time using heuristic functions. The system developed by Delling et al. (41) can support multiple types of optimization objectives for route recommendations in larger map structures, and their methods using overlay topology and parallel computing methods can speed up real-time searches and responses.
@@ 3.2 Probability-based Methods
Generally speaking, large graphs with complete sets of nodes and edges can still be hard to search, even enhanced with heuristics or pruning techniques. Therefore, sampling-based methods have been proposed, which essentially trade off search completeness for probabilistic completeness. In other words, graphs or routes can be sampled in part for construction. One example process is shown in Figure 4. By sampling the passable areas on the map and searching for nearest neighbors, multiple passable paths are generated, and the optimal route is ultimately obtained through a search algorithm.
Figure 4: An example process of probability-based methods. Featured by obtaining the final route through sampling points and performing a proximal search in the map.
For example, Lavalle (54) developed a Probabilistic Road Map (PRM), which constructs a path network in a complex environment by random sampling (transforms the continuous space into a discrete one) and then performs path planning on the sampled network to improve search efficiency in high-dimensional spaces. Besides, the following works, falling in the RRT family, are methods that sample paths. First, Lavalle and Kuffner (55) proposed the Rapidly-exploring Random Tree (RRT) algorithm, which is a single-query algorithm that quickly finds a feasible path from the start point to the end point by random exploring. Its search process resembles a tree growing and expanding in different directions. Kuffner and LaValle (56) later introduced the RRT-Connect algorithm, which can be considered as a bidirectional expansion of two RRT trees. The two trees grow from the start/end point respectively, and they together generate a path when they connect. Karaman and Frazzoli (57) proposed RRT*, which is an asymptotically optimal algorithm. RRT* works similarly to RRT, except that RRT* has a reconnection process when selecting a parent node for a newly added node $v_{new}$. Instead of using the original parent node that generates $v_{new}$, RRT* selects a new parent $v_{pa}$ within a neighbor of $v_{new}$ that has the smallest path cost from $v_{pa}$ to $v_{new}$.
@@
3.3 Biomimetic-based Methods
Biological organisms have inspired many artificial intelligence applications including path planning. Among them, one of the simplest forms of life is called physarum polycephalum, a slime mold that belongs to the amoeba species in the protozoa. This single-celled organism exhibits an interesting phenomenon: on a flat culture medium, its cellular growth path can provide valuable insights for the design of planar transportation networks (123). Therefore, methods that observe, summarize, and apply the rules of biological behaviors in reality to route recommendation are collectively called biomimetic methods. An example process is shown in Figure 5.
Figure 5: An example process of biomimetic-based methods. Featured by obtaining the final route based on the direction and speed of particle or ant colony movement. The particle density near the optimal route is marked high.
@@ 3.4 Clustering-based Methods
Clustering techniques are common useful tools for data dimension reduction and pattern recognition. In our route recommendation setting, clustering is usually used for finding similar route preferences among different users and/or for detecting similar road network structures and traffic flow patterns. By applying appropriate clustering techniques, the original problems can be solved structurally on a smaller scale, thus improving the search efficiency. An example is shown in Figure 6, which is to optimize driving routes based on historical user data.
Figure 6: An example of clustering-based methods. Here clustering is done to provide user density (say, possible taxi customers or historically frequent driving routes). Then an optimal route incorporating user influence is redefined and can differ from the original one.
@@ 3.5 Constraint-based Methods
Constraint-based methods explicitly take into account the route requirements or preferences of users and model them as constraints in the optimization procedure. An example is shown in Figure 7, which is to optimize route search based on user needs imposed as constraints.
Figure 7: An example of constraint-based methods. Featured by establishing final optimization objectives based on user needs.
@@ 4 Deep learning for Route Recommendation
Early machine learning research primarily focused on logical reasoning, knowledge engineering, and simple models with theoretical guarantees. During the last decade, the rapid development of information infrastructures and technologies has allowed much larger-scale computation with big data. Then, more portion of machine learning works are switching to data-driven learning paradigms with more complicated models, where the mainstream is deep learning. Though neural networks have been researched for decades, they are gradually becoming feasible for solving practical problems lately. Compared to classic methods (Section 3), neural networks, especially deep models with delicate structural designs, have huge potential to discover very complex relationships among data. So in this section, we will see why and how deep learning methods prevail in route recommendation research.
@@ 4.1 Why Deep Learning for Route Recommendation?
In the following, we will present deep route recommendation works and categorize them into five types: hybrid models, sequence-based models, graph neural network models, multi-modal models, and deep reinforcement learning models. Hybrid methods usually extend classic methods (such as A*) with neural networks; sequence-based methods mainly model the temporal relations with recurrent neural networks (RNNs); graph neural networks (GNNs) further model spatio-temporal relations; multi-modal methods can process more complex data coming from different modalities; and lastly, deep reinforcement learning methods explicitly focus on making decisions for route recommendations.
DL models are often composed of different modules, for example, using GNNs to extract spatial features, RNNs for time series data, and MLP for feature fusion. As a result, there may be articles appearing in multiple categories. However, we do notice that their novelty contributed to certain parts of the model. So we categorize such articles according to their main contributions and emphasize their novelty under certain categories.
@@ 4.2 Hybrid Approaches
Recall that A* aims to model a cost function $F(x)=g(x)+h(x)$, where $g(x)$ is a base cost function for node $v_{x}$ and $h(x)$ is a heuristic function. The core of A* is to design a good $h(x)$ and estimate both $g(x)$ and $h(x)$ accurately. For example, an early attempt was (85; 90), which proposed using neural networks to automatically learn the cost functions of A* for personalized route recommendation tasks. Wu et al. (86) extended A* using neural networks to model complex traffic information and applied it to the recommendation of fast-traveling routes. Yang et al.(88) proposed a new method named Noah, which combines A* with a graph neural network to calculate an approximate graph editing distance more efficiently and intelligently. Agostinelli et al. (89) proposed and studied batch versions of the A* algorithm. They applied approximate admissible transformations to the heuristic functions that are parameterized by deep neural networks. They showed that these heuristic functions could find optimal or bounded suboptimal solutions. Liu et al. (83) designed a fuel-efficient route recommendation model for long-distance driving based on historical trajectories. They combined neural networks with heuristic algorithms to automatically learn the optimal solutions of A*. Lately, Wang et al. (84) proposed ASNN-FRR, which extracts features from trajectory data with urban road information using AGCRN (124) modules and estimates the values of $g(x)$ and $h(x)$ through a multi-layer perceptron (MLP) network. Besides A*, some works also adopted the classic Dijkstra algorithm in modern models. Jain et al. (32) proposed an inductive algorithm and a generative model called NEUROMLR for robust and reliable route recommendations. They convert the original computational task to the problem of finding the shortest path in a weighted graph and solve it by Dijkstra’s Algorithm, which has guaranteed optimality and accessibility. Liu et al.(13) utilized a deep inverse reinforcement learning (IRL) method to recommend delivery routes. They used the Dijkstra algorithm instead of value iteration to determine the current best strategy and calculate the gradients of the IRL model. Other work such as Huang et al. (87) proposed a multi-task route recommendation framework in contrast to single-planning methods that only work for specific tasks. They used deep learning techniques to extract path representations and employed beam search algorithms to provide recommendations for multiple related tasks.
@@ 4.3 Sequence-based Approaches
The following examples are RNN-based methods for route recommendation. An early work (93) used RNN to extract feature representations of different POI nodes and added maximum difference constraints to recommend a path with diverse scenes for users. Wang et al. (85; 90) proposed a personalized route recommendation method, which uses GRU segments to extract features from numerous user trajectories and employs the attention mechanism to increase the prediction accuracy. Similarly, Wen et al. (31) used GRUs to extract the temporal correlations in trajectory data, which helped to improve route recommendations for cargo deliveries. Huang et al. (87) proposed a multi-task prediction framework for route recommendation, which uses LSTM to fuse POI features, user features, and current path features, and applies a greedy network search algorithm to output results for multiple tasks. Hahn et al. (91) aimed to monitor collision information on roads and then recommend safer routes. They extracted features of collision accidents using LSTM and used them to predict the next step of moving positions. Liu et al. (83) combined a bidirectional GRU and attention mechanism to estimate the fuel consumption in long-distance driving routes and recommended energy-efficient routes for drivers. Fu and Lee (92) designed a progressive route recommendation method based on generative networks, which extracts knowledge from the selection of the next node using LSTM. Wang et al. (84) introduced a travel route recommendation algorithm using AGCRN modules (124), which combines an RNN and a graph convolutional neural network (GCN) to capture spatial and temporal correlations in a fine-grained dimensional space.
Besides RNNs which are autoregressive models, the Transformers (128) are attention-based models that can process much longer sequences without suffering from the forgetting problem. The Transformer architecture consists of several key modules such as positional encoding, self-attention (Eq. (11)), multi-head attention (Eq. (12)), and feed-forward neural networks.
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (11)
Where matrices $Q$, $K$, and $V$ correspond to the query, key, and value components, respectively, and $d_{k}$ denotes the dimension of the key.
$\text{MultiHead}(Q,K,V)=\text{Concat}(head_{1},...,head_{h})W^{O}$ (12)
where $\text{where }head_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$
Where matrices $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, and $W^{O}$ are learnable weight matrices within the model. So we see that Transformers essentially differ from autoregressive models by using attention to process long sequences.
@@
Therefore, many studies began to explore route recommendation methods based on Transformer. For instance, Bhumika and Debasis Das (95) introduced a multi-task learning framework, which uses Transformers to extract spatial, temporal, and semantic correlations, and adopted epsilon constraint techniques to optimize routes that satisfy multiple objective functions. Wang et al. (96) studied long-distance travel trajectories, from which they extracted fuel-consumption factors using Transformer. Further combining a genetic optimization algorithm, they developed a model that could recommend more fuel-economic routes. Wang et al. (97) modeled route recommendation as a process of inferring a high-sampling rate fine trajectory from a low-sampling rate coarse trajectory composed of origin-destinations and way-points. Their proposed Seq2Seq model using multi-head self-attention mechanisms can automatically adjust the sampling weights to capture more accurate spatio-temporal correlation. Wang et al. (98) proposed a new de-biased representation learning method named TripRec to implement the sequence generation method from Query to Trip. Based on the encoded representations, they further used Transformer for decoding and generated final routes. Ngai Lam Ho and Kwan Hui Lim (99) proposed POIBERT, which uses the BERT language model to recommend personalized itineraries based on users’ preferences and their past POI selections.